

********************************************************************
UARS WEEKLY STATUS REPORT - 02 APR 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE was operating during daylight only for four days this week in
support of the yaw maneuever. Otherwise, SOLSTICE continued to operate
in good health gathering solar data on all available orbital tracking
periods. On average, we received 62% of each day's data amounting to
about 400 minutes of solar observing time per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg II,
occultation, solar pointing offset and high resolution continuous
monitoring scans for periods of UARS real-time telemetry or onboard
tape recording.
As a result of the 3-TDRS operation, we have been able to fully resume
nearly all of the scans that were routinely at the start of the UARS
mission. The weekly high (0.15 nm) resolution full spectrum
(115-410 nm) solar scans, the daily low resolution full spectrum
scans, and the daily scans of selected solar features at high
resolution (aka continuous monitoring) will now be performed on their
respective cadences without data gaps resulting from the limitations
of available telemetry. Resumption of the remaining activities done
earlier in the UARS mission (ozone occulatations and gridded offset
manuevers) are now being considered.
Again, many thanks to those who made the 3-TDRS operation possible.
***********
PEM
***********
Level 2 PEM HEPSA data files are being reprocessed for submission to the
DAAC. This data has been designated as version 2; however, there is no format
difference from the version 1 data format. Version 2 HEPSA data is being
generated because it was discovered that there exists processing errors within
some of the HEPSA version 1 data files. Thus, if you are using any HEPSA
version 1 data files from the DAAC, please replace them with the version 2
data files. Currently, HEPSA data between 1991274 and 2003365 has been
reprocessed and archived at the GSFC DAAC.
Level 2 PEM MEPS data file documentation is complete and submitted to
the DAAC. MEPS data files are version 1. Processed MEPS data files at the
DAAC archive are between 1991274 and 1995299. MEPS data files between
1995300 and 1995365 are waiting for DAAC collection.
PEM continues to receive NMC data files from NOAA. No problems have
been encountered. NOAA has changed the binary structure of the NMC data and
PEM is now receiving two data streams, one in the original VAX format and the
other from their new binary format.
The PEM data from the GUVI conjunction periods may be found at
http://wwwpem.space.swri.edu/guvi-pem.html. The additional data for the
October and November storms of 2003 has been moved and is now located at
http://climatology.space.swri.edu/uars_fromRudy.html. For those users of
SDDAS, update your UARS Meta data and these data times will appear
with full functionality.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 20 March 2004 (Orbit 68518, GMT Day 080) through 26
March 2004 (Orbit 68623, GMT Day 086).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 086/0143 68609 TO LOW POWER MODE
HALOE 086/0806 68613 Power OFF due to low power
The beta angle ranged between the angles of 40.6 degrees to 12.1
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 29 March 2004 (DOY 089).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.92 to 2.54 degrees C. Additional
battery performance data for each battery on 19 March and 26 March
2004 are provided in Appendix D.
The clock error ranged from 05.6 to -11.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 01
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 02
HRDI = 01
HALOE = 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT transitioned the CMS to the new software on 15 March 2004.
All loads are being built on this new version without incident. This
was done to prepare for the start of 3-TDRS operations, which began on
17 March 2004. The FOT is now taking 35-38 events per day (up from 26-
28 per day under 2-TDRS operations) and now collecting data for an
additional 18-20% more of each orbit. This 3-TDRS mode will be the
nominal mode from now on. This new mode will be closely examined to
see if any slight modifications may be needed in the future, but none
are planned at this time. The FOT would like to once again thank all
those involved in making this happen and thank science for being
patient with this effort. Any feedback from science on how this mode
is working for them would be greatly appreciated.
MEETINGS
-None this report period
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 29 March 2004 (DOY 089). This
maneuver will be a Reverse to Forward Flight maneuver (Southern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 26 March 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68605 085/2008 04-089 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68607 085/2147 04-090 OBC/CLK CLK ADJ FR 20.6 TO -20.6 MSEC/DAY
68609 086/0143 04-091 HRDI HRDI TO LOW POWER MODE
68613 086/0806 04-092 HALOE POWER OFF DUE TO LOW POWER
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 26 March 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
OBSBATVEON3 Y-LO Battery 3 Voltage at EON 68613
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 26 March 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68605 085/2012 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
68607 085/2152 CLK ADJ FR 20.6 TO -20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 26 March 2004
19 March 2004 - GMT Day 079 Beta = 40.7 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-095.2 +5.79/+4.50 +27.2/-15.2 24.8 1.04 21.5
3 +22.4/-078.4 +4.82/+3.55 +26.4/-15.2 24.6 1.04 21.0
26 March 2004 - GMT Day 086 Beta = 12.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.63 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-134.4 +5.79/+3.87 +28.0/-14.4 23.7 1.02 19.8
3 +22.4/-089.6 +5.14/+2.60 +27.2/-14.4 23.5 1.02 19.5
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None This Report Period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 26 March 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT 68524 080/1004 30832 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-015 FOT 68430 074/0401 35722 Late Acquisition
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-014 FOT See List Mar 04 30832 Generic Late Acquisition